# GenAINet: Enabling Wireless Collective Intelligence via Knowledge Transfer and Reasoning

### v2

@@ V Challenges and Opportunities
We envision GenAINet as a vital part for enabling collective intelligence in 6G. However, there are several challenges which opens up new research opportunities.
@@
LLMs are auto-regressive generative models pre-trained on natural text to predict the next token. Therefore LLMs hardly predict high-level semantics, which introduces high information redundancy and computation cost in training and inference. Besides, LLMs’ generalization capability is usually limited in domains not included in the training data. RAG and fine-tuning are commonly used to enhance LLMs’ knowledge while being expensive and inflexible. Furthermore, grounding LLMs’ knowledge to real world representation is challenging. Reinforcement learning is studied to optimize LLM online, which is not suitable for computational limited devices.
@@
World model is a hierarchical, modular model predicting future representations of the state of the world [14]. It is trained to predict the high-level abstract instead of raw data. For example, I-JEPA [15] is trained to predict missing embedding on an image, which shows less training data and computational effort. Furthermore, hierarchical JEPA (H-JEPA) allows to learn higher level abstract representation which is effective for long-term prediction and eliminate the irrelevant details. Built on H-JEPA, hierarchical planning can be done by predicting state transition on abstract space, which could handle uncertain environment with minimal cost. JEPA framework gives promises for deploying world model on wireless devices than LLMs. Since RF propagation is more abstract than text, training RF-JEPA could be more efficient than RF-GPT. However, JEPA on multi-modality and planning still require further research.
@@
While GenAI agents are used in various domains, applying them in wireless networks is still challenging. First, a network has a complex hierarchical architecture from RF to service layers, making LLMs difficult to decompose tasks for every network element. Second, orchestrating numerous agents’ behavior is difficult for networks deployed in large scale geographical areas. Therefore, current LLMs are used only in small domains, mostly service layer. Moreover, future networks require much higher reliability and robustness, where the uncertainty of GenAI models is problematic. Finally embedding RF signals into LLMs is challenging, primarily due to the unavailability of appropriate large datasets and the inherent nature of RF signals, which is both spectral and spatial. This fundamentally differs from the textual data.
@@ VI Conclusion
In this paper, we introduced the GenAINet, a knowledge driven communication and reasoning network as a promising enabler for collective intelligence. The architecture of GenAINet leverages unified LLM-powered GenAI agents to optimize both network protocols and applications. Our semantic-native framework can learn structure of semantic concept or abstract from multi-modal raw data, and build semantic knowledgebase to achieve effective communication and reasoning. Agents utilize LLMs with minimal semantic concepts retrieved locally and remotely through semantic communications to effectively plan reasoning paths and produce effective decisions for a task. We investigated a use case of transferring semantic knowledge between teacher and student agents to improve on-device query with less communication costs, as well as an example of distributed agents communicate with reasoning to improve wireless power control. We demonstrated that our multi-agent GenAINet can unleash the power of collective intelligence, and addressed that developing a semantic-native Telecom world model is an essential path towards network collective intelligence.
Biographies
Hang Zou (hang.zou@tii.ae) is a Researcher at Technology Innovation Institute, UAE.
Qiyang Zhao (qiyang.zhao@tii.ae) is a Principal Researcher at Technology Innovation Institute, UAE.
Lina Bariah (lina.bariah@ieee.org) is an Adjunct Professor at Khalifa University, UAE.
Yu Tian (yu.tian@tii.ae) is a Researcher at Technology Innovation Institute, UAE.
Mehdi Bennis (mehdi.bennis@oulu.fi) is a Professor at University of Oulu, Finland.
Samson Lasaulce (samson.lasaulce@ku.ac.ae) is a Chief Research Scientist with Khalifa University (UAE) and a CNRS Director of Research (France).
Mérouane Debbah (merouane.debbah@ku.ac.ae) is a Professor at Khalifa University and a Senior Advisor at Technology Innovation Institute, UAE.
Faouzi Bader (carlos-faouzi.bader@tii.ae) is the Senior Director of Telecom Unit at Technology Innovation Institute, UAE.

### v3

@@ V Challenges and Opportunities
Number of Transmitters Absolute Rate Gap (kbps) Total Power (W) Messages/Txs (GenAINet)
DPC GenAI Alone GenAINet DPC GenAI Alone GenAINet
2 1.747 1.625 1.397 16.536 8.842 10.700 3.72
4 2.086 1.967 1.736 28.254 17.156 15.974 5.00
10 2.433 2.403 2.249 76.785 34.696 30.622 3.80
TABLE II: Performance comparison in terms of absolute rate gap, total power, and messages per transmitter (GenAINet only) for DPC, GenAI agents alone with memory and reasoning, and GenAINet allowing cooperation through natural language between agents.
@@
We envision GenAINet as a vital part for enabling CI in 6G. However, there are several challenges to integrate LLMs in wireless networks which opens up new research opportunities.
Edge LLMs The vast majority of existing LLMs are deployed in the cloud, limiting their applicability on devices due to constraints pertinent to bandwidth, latency, and security. Despite their promising capabilities, deploying mainstream LLMs on edge devices is extremely challenging. First, deploying large-scale LLMs would be expensive for ordinary devices. Additionally, LLMs have long inference time when solving sophisticated tasks, which significantly increases the communication latency. Finally, standards and protocols of wireless networks renew consistently, leading knowledge synchronization and updating of LLMs energy-consuming through conventional techniques. Therefore it is crucial to develop efficient techniques to adapt LLMs to wireless environment. Model compression, inference acceleration and selective knowledge sharing will mitigate these challenge potentially. Model compression in particular allows significant reductions in terms of required memory to be achieved. Model compression relies on key techniques such as quantization [36], pruning [37], distillation [38], low-rank factorization [39], parameter sharing [40], weight clustering [41], sparse attention mechanisms [42], or dynamic inference techniques [43]. For instance, for quantization, significant gains can be obtained by quantizing float or double numbers (32 and 64 bits respectively) which are used to describe the LLM parameters. In the BitNet paper [44], it is shown to quantize parameters with only 1.58 bit and obtain an energy consumption which is decreased by a factor of 71.4x. Also, efficient attention mechanism optimization techniques such as FlashAttention [45] and efficient inference techniques including PageAttention [46], KV caching and prefill-decode separation will lead to a considerable reduction of inference time. Finally task/goal-oriented compression techniques helps to further reduce the communication overhead between agents [47].
@@
Efficient Adaptation LLMs are auto-regressive generative models pre-trained on natural text to predict the next token. Therefore LLMs hardly predict high-level semantics, which introduces high information redundancy and computation cost in training and inference. Besides, LLMs’ generalization capability is usually limited in domains not included in the training data. RAG and fine-tuning are commonly used to enhance LLMs’ knowledge while being expensive and inflexible. Furthermore, grounding LLMs’ knowledge to real world representation is challenging. Reinforcement learning is studied to optimize LLM online, which is not suitable for computational limited devices.
@@
World Model is a hierarchical, modular model predicting future representations of the state of the world [48]. It is trained to predict the high-level abstract instead of raw data. For example, I-JEPA [49] is trained to predict missing embedding on an image, which shows less training data and computational effort. Furthermore, hierarchical JEPA (H-JEPA) allows to learn higher level abstract representation which is effective for long-term prediction and eliminate the irrelevant details. Built on H-JEPA, hierarchical planning can be done by predicting state transition on abstract space, which could handle uncertain environment with minimal cost. JEPA framework gives promises for deploying world model on wireless devices than LLMs. Since RF propagation is more abstract than text, training RF-JEPA could be more efficient than RF-GPT. However, JEPA on multi-modality and planning to enable CI still require further research.
@@
RF Foundation Models While GenAI agents are used in various domains, applying them in wireless networks is still challenging. First, a network has a complex hierarchical architecture from RF to service layers, making LLMs difficult to decompose tasks for every network element. Second, orchestrating numerous agents’ behavior is difficult for networks deployed in large scale geographical areas. Therefore, current LLMs are used only in small domains, mostly service layer. Moreover, future networks require much higher reliability and robustness, where the uncertainty of LLMs is problematic. Finally embedding RF signals into LLMs is challenging, primarily due to the unavailability of appropriate large datasets and the inherent nature of RF signals, which is both spectral and spatial. This fundamentally differs from the textual data.
@@ VI Conclusion
In this paper, we introduced the GenAINet framework, a knowledge-driven communication and reasoning network, as a promising enabler for collective intelligence. The architecture of GenAINet leverages unified LLM-powered GenAI agents to optimize both network protocols and applications. Agents in our semantic-native framework can extract unified semantics from heterogeneous raw data in the environment or from other agents, and leverage a semantic knowledge model to achieve effective communication and reasoning. Agents utilize LLMs with semantics retrieved locally and remotely through semantic communications to effectively plan reasoning paths and make informed decisions for a task. We investigated a use case involving the compression and transfer of common knowledge between two agents to improve on-device query performance with reduced communication costs, as well as an example where distributed agents communicate using reasoning to accomplish general wireless power control problems without predefined communication protocols. We demonstrated that our multi-agent GenAINet can unleash the power of collective intelligence with flexible levels of collaboration, much like human interactions, and highlighted some fundamental challenges and future directions for leveraging LLMs in 6G networks to enable collective intelligence. Of course, challenges remain, such as making the proposed framework scalable. In particular, having sufficiently small and fast LLMs is necessary for their massive deployment in wireless networks. Moreover, as both a fundamental and practical challenge, it is imperative to design LLMs that allow some guarantees in terms of QoS and efficiently manage energy consumption.